# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### 06 Mar 2026 v2

@@ 1.1 Formulation of the EVP model
Sea-ice dynamics has a strong influence on the global climate, for instance, through heat exchange with the atmosphere and the ocean around the Arctic area as well as the reflection of sunlight. Despite the development of sea-ice modelling over many decades, its complexity continues to pose significant challenges for numerical simulations of the sea-ice cover (cf. Ref. [29]). For these reasons, an enhanced theoretical understanding of the fundamental dynamical sea-ice models and their rigorous mathematical analysis is of great importance and substantial interest.
@@
The EVP model was introduced in Ref. [39] as a numerical regularisation of the Hibler model from Ref. [36]. The latter model describes sea-ice dynamics by using a viscous-plastic rheology, and the viscosity becomes singular for small strain rates and degenerate for large strain rates.
@@
$D(u)\coloneqq\frac{1}{2}[\nabla u+(\nabla u)^{\top}].$ (1)
@@
$\overline{\mathcal{D}}\coloneqq\biggl(\frac{2}{\overline{e}^{2}}\bigg|D(u)-\frac{1}{2}[\operatorname{Tr}D(u)]\mathbb{I}_{2}\bigg|^{2}+\big|\operatorname{Tr}D(u)\big|^{2}\biggr)^{1/2}.$ (2)
@@
Then, the momentum equations of the Hibler sea-ice model can be written as[36]
$m\partial_{t}u+mu\cdot\nabla u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+m\Omega u^{\perp}-mg\nabla H_{0},$ (3a)
@@
$\dfrac{\overline{e}^{2}\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u).$ (3b)
@@
Here $\sigma:\mathbb{T}^{2}\times[0,T]\rightarrow\mathbb{R}^{2\times 2}$ is the stress tensor, and $P>0,m>0$, and $\Omega$ are given constants which denote the pressure, the mass per unit area and the coefficient of the Coriolis force, respectively. The functions $\mathcal{T}_{a}$ and $\mathcal{T}_{w}$ denote the wind and ocean stresses, and are defined in equations (3d) and (3e) below, respectively. The given function $H_{0}$ denotes the ocean surface topography. We also recall the notation $u^{\perp}=(-u_{2},u_{1})^{\top}$. In order to fix ideas, we have omitted the thermodynamical equations (i.e. we assume the mean ice thickness $h$ and the ice compactness $A$ to be constant) and purely focus on the dynamical part of the EVP model (as is also done for the EVP model as originally introduced in Ref. [39]). Furthermore, we assume that $P$ and $m$ are given constants (as we have taken $h$ and $A$ to be constant). The analysis of the EVP model coupled with the thermodynamical equations is a subject of future work.
@@
One can easily check from (3b) that $\sigma$ becomes singular as the strain rate $\overline{\mathcal{D}}$ (see (2)) approaches zero, and becomes degenerate as $\overline{\mathcal{D}}$ approaches infinity. This special structure, in particular, causes significant difficulties for both the analysis as well as the numerical simulations. In fact, in Hibler’s original paper, a cutoff of $\overline{\mathcal{D}}$ is introduced to avoid this very problem.[36]
@@
In order to facilitate the usage of explicit numerical schemes and therefore also of parallel computing, it was proposed in Ref. [39] to modify the viscous-plastic sea-ice rheology (3b) by adding an artificial elastic component. That is, one replaces the diagnostic, constitutive relation (3b) by the prognostic, dynamically relaxed equation for the stress tensor
@@
$\dfrac{1}{\mathcal{E}}\partial_{t}\sigma+\dfrac{\overline{e}^{2}\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (3c)
@@
where $\mathcal{E}>0$ is the (artificial) elastic modulus, which is constant. Formally, by sending $\mathcal{E}\rightarrow\infty$ in (3c), one can recover (3b). In addition, we will show in this paper, that a cutoff of the strain rate $\overline{\mathcal{D}}$ is not required in the dynamical case with Eq. (3c) at hand for $\mathcal{E}>0$. Note that in contrast to that of the Hibler model, the stress tensor of the EVP model has become an independent dynamical quantity, which requires its own initial condition which need to be chosen adequately in numerical simulations. Moreover, the parabolic structure of the full Hibler model (3a)-(3b) is lost due to this modification, although it reemerges as a damping term in the EVP model.
@@
The EVP model, which is given by equations (3a) and (3c), is one of the most important dynamical sea-ice models and is a fundamental component of many climate models,[7, 29, 44] including the Los Alamos sea-ice model (CICE),[51] the Louvain-la-Neuve (LIM2) sea-ice model,[7] the Massachusetts Institute of Technology general circulation model (MITgcm),[60] and the Finite-Volume sea-ice–ocean model (FESOM)[25] (an overview of several models can be found in Ref. [37]). Furthermore, the wind stress $\mathcal{T}_{a}$ and the ocean stress $\mathcal{T}_{w}$ are defined as follows
$\mathcal{T}_{a}$ $\coloneqq c_{a}\rho_{a}\lvert U_{a}\rvert\bigg(U_{a}\cos\phi+U_{a}^{\perp}\sin\phi\bigg),$ (3d)
$\mathcal{T}_{w}$ $\coloneqq c_{w}\rho_{w}\lvert U_{w}-u\rvert\bigg[(U_{w}-u)\cos\theta+(U_{w}-u)^{\perp}\sin\theta\bigg],$ (3e)
@@
respectively. The other notation is introduced in Table 1. For the purposes of this paper, the geostrophic wind and ocean currents $U_{a}$ and $U_{w}$ can be arbitrary functions in $C([0,T];H^{4}(\mathbb{T}^{2}))$, but we note that the ocean velocity $U_{w}$ is typically chosen as follows, cf. e.g. Refs. [25, 38],
$U_{w}=\begin{pmatrix}\frac{0.1(2y-L)}{L}\\ \frac{-0.1(2x-L)}{L}\end{pmatrix}$, (4)
@@
where $L$ is the length scale of the domain $0,L]^{2}$. In this paper we will take $L=1$. Note that this example is not a spatially periodic vector field. Similarly, a typical wind velocity is given by (cf. Ref. [38])
$U_{a}=\begin{pmatrix}5+\left[\sin\left(\frac{2\pi t}{T}\right)-3\right]\sin\left(\frac{2\pi x}{L}\right)\sin\left(\frac{\pi y}{L}\right)\\ 5+\left[\sin\left(\frac{2\pi t}{T}\right)-3\right]\sin\left(\frac{2\pi y}{L}\right)\sin\left(\frac{\pi x}{L}\right)\end{pmatrix}$. (5)
@@
The goal of this paper is to establish the global well-posedness of a regularised version of the EVP system, i.e. equations (3a) and (3c)-(3e). To simplify our presentation, we will fix $\overline{e}\equiv 2$ in the rest of this paper, as is commonly done in the sea-ice modelling literature. However, we emphasise that the results of this paper do not depend on the choice of $\overline{e}$.
@@
$U_{a}$ geostrophic wind Eq. (5)
$U_{w}$ geostrophic ocean current Eq. (4)
Table 1: Index of notation and typical values of several of the parameters, which are taken from Refs. [39, 38, 50, 61].
@@ 1.2 The linear ill-posedness of the unregularised EVP model
$\partial_{t}u=\partial_{x}\sigma,$ (6a)
$\partial_{t}\sigma+\frac{5\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}}{2P}\sigma+\frac{\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}}{2}=\partial_{x}u,$ (6b)
where $u$ and $\sigma$ are now scalar quantities, and we have regularised and simplified the strain rate $\overline{\mathcal{D}}=\sqrt{\frac{5}{4}}\lvert\partial_{x}u\rvert$ from (2) by $\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}$ for some $\epsilon>0$. Note that we have omitted the external forces as well as the advection term (cf. point $(ii)$ of Remark 1.3). By linearising system (6) around a given solution $(\overline{u},\overline{\sigma})$ to (6) we obtain the following system of equations
$\partial_{t}u_{l}=\partial_{x}\sigma_{l},$ (7a)
$\partial_{t}\sigma_{l}=-\frac{5}{2P}\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}\sigma_{l}+\bigg[1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}\bigg]\partial_{x}u_{l},$ (7b)
where $(u_{l},\sigma_{l})$ is the solution of the linearised system (7a)-(7b). Then, after taking a spatial derivative of Eq. (7b) and using Eq. (7a), one has
$\partial_{tt}u_{l}=$ $-\frac{5}{2P}\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}\partial_{t}u_{l}+\bigg[1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}\bigg]\partial_{xx}u_{l}$ (8)
@@
$1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}<0,$ (9)
then Eq. (8) becomes locally elliptic near the point $(x_{0},t_{0})$, which leads to an instability and therefore the ill-posedness in the sense of Hadamard.[34] In a forthcoming paper Ref. [8], we will address the linear ill-posedness of the (unregularised) EVP model in all Sobolev spaces in a rigorous manner using this new instability (ill-posedness) that we have discussed above.
@@
We emphasise that the instability described above is fundamentally different than the (formal) ill-posedness/instability results that have been reported in the sea-ice modelling/physics literature.[30, 31, 32, 65] These results rely crucially on the absence of the viscosity cutoff (i.e. $\epsilon=0$) and are found mainly for the Hibler model (with the exception of Refs. [56, 71], which consider the (revised) EVP model as well). The rigorous analysis in Ref. [57] demonstrates that the linear ill-posedness is not present in the case of the Hibler model with $\epsilon>0$. The linear instability of the EVP model described in this section is present regardless of the value of $\epsilon$ and therefore pertains to a new kind of instability that has not been reported in the literature before, and in particular it does not apply to the Hibler model.
@@ 1.3 The main result of this paper
We consider the following regularisation of the EVP model (i.e. equations (3a) and (3c)):
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (10a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (10b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (10c)
In this model we have included the Voigt regularisation on the elastic part of the rheology in the EVP model[39] by means of the inclusion of the term $\alpha^{2}\partial_{t}\Delta\sigma$ for some fixed $\alpha>0$ in Eq. (10b). The drag forces $\mathcal{T}_{a}$ and $\mathcal{T}_{w}$ are taken as in Eqs. (3d) and (3e), respectively. Moreover, for the sake of simplicity we have chosen $m\equiv 1$. Note that throughout this paper we will assume that $\sigma(x,t)$ is a symmetric matrix, and in Lemma 2.4 we show rigorously that the stress tensor $\sigma$ remains symmetric if $\sigma_{0}$, i.e. initially, is symmetric. Next, we will simplify the strain rate $\overline{\mathcal{D}}$ from Eq. (2) by replacing it with
$\mathcal{D}\coloneqq\lvert D(u)\rvert$, (11)
@@
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (12a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (12b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (12c)
@@
Let $u_{0}\in H^{2}(\mathbb{T}^{2})$ and $\sigma_{0}\in H^{3}(\mathbb{T}^{2})$ (which is symmetric) be given. Moreover, we assume that $P>0$ is a given constant, $U_{a},U_{w},H_{0}\in C([0,T];H^{4}(\mathbb{T}^{2}))$ and we let $\theta\in\left[0,\frac{\pi}{4}\right]$. The Voigt-EVP system (12) has a unique global strong solution $(u,\sigma)$ on the time interval $[0,T]$ for any $T>0$. The solution $(u,\sigma)$ enjoys the following regularity
@@
    The Voigt regularisation in Eq. (12b) is in the same spirit as the original numerical regularisation of the Hibler model from Ref. [39]. Namely, we seek an inviscid regularisation of the EVP model, which is suitable for numerical computations. The advantage of the Voigt regularisation is that on the one hand, like the original EVP model, it shares the same steady state(s) with the Hibler model. On the other hand, the Voigt regularisation constitutes a relatively modest (nonviscous) modification of the (artificial) elastic component of the rheology in the EVP model.
@@
    For the momentum equation of the Voigt-EVP model, we have omitted the advection term, in which we follow the approach in Ref. [39]. As it has been remarked in Ref. [39], the advection term is at least one order of magnitude smaller than the acceleration term and therefore it has not been included in the EVP model. In the evolution equation for the stress tensor we also do not include an advection term, as it would affect the steady states of the EVP model, which would then lose its purpose as an approximation of the Hibler model.
@@
    We note that the assumption $\theta\in\left[0,\frac{\pi}{4}\right]$ on the water turning angle, lies within the standard range in sea-ice modelling (see for example Refs. [35, 53]).

### 20 Mar 2026 v3

@@ 1.1 Formulation of the EVP model
Sea-ice dynamics has a strong influence on the global climate, for instance, through heat exchange with the atmosphere and the ocean around the Arctic area as well as the reflection of sunlight. Despite the development of sea-ice modelling over many decades, its complexity continues to pose significant challenges for numerical simulations of the sea-ice cover (cf. [29]). For these reasons, an enhanced theoretical understanding of the fundamental dynamical sea-ice models and their rigorous mathematical analysis is of great importance and substantial interest.
@@
The EVP model was introduced in [38] as a numerical regularisation of the Hibler model from [36]. The latter model describes sea-ice dynamics by using a viscous-plastic rheology, and the viscosity becomes singular for small strain rates and degenerate for large strain rates.
@@
$D(u)\coloneqq\frac{1}{2}[\nabla u+(\nabla u)^{\top}].$ (1.1)
@@
$\overline{\mathcal{D}}\coloneqq\biggl(\frac{2}{\overline{e}^{2}}\bigg|D(u)-\frac{1}{2}[\operatorname{Tr}D(u)]\mathbb{I}_{2}\bigg|^{2}+\big|\operatorname{Tr}D(u)\big|^{2}\biggr)^{1/2}.$ (1.2)
@@
Then, the momentum equations of the Hibler sea-ice model [36] can be written as
$m\partial_{t}u+mu\cdot\nabla u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+m\Omega u^{\perp}-mg\nabla H_{0},$ (1.3a)
@@
$\dfrac{\overline{e}^{2}\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u).$ (1.3b)
@@
Here $\sigma:\mathbb{T}^{2}\times[0,T]\rightarrow\mathbb{R}^{2\times 2}$ is the stress tensor, and $P>0,m>0$, and $\Omega$ are given constants which denote the pressure, the mass per unit area and the coefficient of the Coriolis force, respectively. The functions $\mathcal{T}_{a}$ and $\mathcal{T}_{w}$ denote the wind and ocean stresses, and are defined in equations (1.3d) and (1.3e) below, respectively. The given function $H_{0}$ denotes the ocean surface topography. We also recall the notation $u^{\perp}=(-u_{2},u_{1})^{\top}$. In order to fix ideas, we have omitted the thermodynamical equations (i.e. we assume the mean ice thickness $h$ and the ice compactness $A$ to be constant) and purely focus on the dynamical part of the EVP model (as is also done for the EVP model as originally introduced in the paper [38]). Furthermore, we assume that $P$ and $m$ are given constants (as we have taken $h$ and $A$ to be constant). The analysis of the EVP model coupled with the thermodynamical equations is a subject of future work.
@@
One can easily check from (1.3b) that $\sigma$ becomes singular as the strain rate $\overline{\mathcal{D}}$ (see (1.2)) approaches zero, and becomes degenerate as $\overline{\mathcal{D}}$ approaches infinity. This special structure, in particular, causes significant difficulties for both the analysis as well as the numerical simulations. In fact, in Hibler’s original paper [36], a cutoff of $\overline{\mathcal{D}}$ is introduced to avoid this very problem.
@@
In order to facilitate the usage of explicit numerical schemes and therefore also of parallel computing, it was proposed in [38] to modify the viscous-plastic sea-ice rheology (1.3b) by adding an artificial elastic component. That is, one replaces the diagnostic, constitutive equation (1.3b) by the prognostic, dynamically relaxed equation for the stress tensor
@@
$\dfrac{1}{\mathcal{E}}\partial_{t}\sigma+\dfrac{\overline{e}^{2}\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (1.3c)
@@
where $\mathcal{E}>0$ is the (artificial) elastic modulus, which is constant. Formally, by sending $\mathcal{E}\rightarrow\infty$ in (1.3c), one can recover (1.3b). In addition, we will show in this paper, that a cutoff of the strain rate $\overline{\mathcal{D}}$ is not required in the dynamical case with equation (1.3c) at hand for $\mathcal{E}>0$. Note that in contrast to that of the Hibler model, the stress tensor of the EVP model has become an independent dynamical quantity, which requires its own initial condition which need to be chosen adequately in numerical simulations. Moreover, the parabolic structure of the full Hibler model (1.3a)-(1.3b) is lost due to this modification, although it reemerges as a damping term in the EVP model.
@@
The EVP model, which is given by equations (1.3a) and (1.3c), is one of the most important dynamical sea-ice models and is a fundamental component of many climate models [7, 29, 44], including the Los Alamos sea-ice model (CICE) [51], the Louvain-la-Neuve (LIM2) sea-ice model [7], the Massachusetts Institute of Technology general circulation model (MITgcm) [60], and the Finite-Volume sea-ice–ocean model (FESOM) [25] (an overview of several models can be found in [37]). Furthermore, the wind stress $\mathcal{T}_{a}$ and the ocean stress $\mathcal{T}_{w}$ are defined as follows
$\mathcal{T}_{a}$ $\coloneqq c_{a}\rho_{a}\lvert U_{a}\rvert\bigg(U_{a}\cos\phi+U_{a}^{\perp}\sin\phi\bigg),$ (1.3d)
$\mathcal{T}_{w}$ $\coloneqq c_{w}\rho_{w}\lvert U_{w}-u\rvert\bigg[(U_{w}-u)\cos\theta+(U_{w}-u)^{\perp}\sin\theta\bigg],$ (1.3e)
@@
respectively. The other notation is introduced in Table 1. For the purposes of this paper, the geostrophic wind and ocean currents $U_{a}$ and $U_{w}$ can be arbitrary functions in $C([0,T];H^{4}(\mathbb{T}^{2}))$, but we note that the ocean velocity $U_{w}$ is typically chosen as follows, cf. e.g. [40, 25],
$U_{w}=\begin{pmatrix}\frac{0.1(2y-L)}{L}\\ \frac{-0.1(2x-L)}{L}\end{pmatrix}$, (1.4)
@@
where $L$ is the length scale of the domain $0,L]^{2}$. In this paper we will take $L=1$. Note that this example is not a spatially periodic vector field. Similarly, a typical wind velocity is given by (cf. [40])
$U_{a}=\begin{pmatrix}5+\left[\sin\left(\frac{2\pi t}{T}\right)-3\right]\sin\left(\frac{2\pi x}{L}\right)\sin\left(\frac{\pi y}{L}\right)\\ 5+\left[\sin\left(\frac{2\pi t}{T}\right)-3\right]\sin\left(\frac{2\pi y}{L}\right)\sin\left(\frac{\pi x}{L}\right)\end{pmatrix}$. (1.5)
@@
The goal of this paper is to establish the global well-posedness of a regularised version of the EVP system, i.e. equations (1.3a) and (1.3c)-(1.3e). To simplify our presentation, we will fix $\overline{e}\equiv 2$ in the rest of this paper, as is commonly done in the sea-ice modelling literature. However, we emphasise that the results of this paper do not depend on the choice of $\overline{e}$.
@@
$U_{a}$ geostrophic wind equation (1.5)
$U_{w}$ geostrophic ocean current equation (1.4)
Table 1: Index of notation and typical values of several of the parameters, which are taken from [38, 40, 61, 50].
@@ 1.2 The linear ill-posedness of the unregularised EVP model
$\partial_{t}u=\partial_{x}\sigma,$ (1.6a)
$\partial_{t}\sigma+\frac{5\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}}{2P}\sigma+\frac{\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}}{2}=\partial_{x}u,$ (1.6b)
where $u$ and $\sigma$ are now scalar quantities, and we have regularised and simplified the strain rate $\overline{\mathcal{D}}=\sqrt{\frac{5}{4}}\lvert\partial_{x}u\rvert$ from (1.2) by $\sqrt{\lvert\partial_{x}u\rvert^{2}+\epsilon^{2}}$ for some $\epsilon>0$. Note that we have omitted the external forces as well as the advection term (cf. point $(ii)$ of Remark 1.3). By linearising system (1.6) around a given solution $(\overline{u},\overline{\sigma})$ to (1.6) we obtain the following system of equations
$\partial_{t}u_{l}=\partial_{x}\sigma_{l},$ (1.7a)
$\partial_{t}\sigma_{l}=-\frac{5}{2P}\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}\sigma_{l}+\bigg[1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}\bigg]\partial_{x}u_{l},$ (1.7b)
where $(u_{l},\sigma_{l})$ is the solution of the linearised system (1.7a)-(1.7b). Then, after taking a spatial derivative of equation (1.7b) and using equation (1.7a), one has
$\partial_{tt}u_{l}=$ $-\frac{5}{2P}\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}\partial_{t}u_{l}+\bigg[1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}\bigg]\partial_{xx}u_{l}$ (1.8)
@@
$1-\frac{5}{2P}\frac{\overline{\sigma}\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}-\frac{1}{2}\frac{\partial_{x}\overline{u}}{\sqrt{\lvert\partial_{x}\overline{u}\rvert^{2}+\epsilon^{2}}}<0,$ (1.9)
then equation (1.8) becomes locally elliptic near the point $(x_{0},t_{0})$, which leads to an instability and therefore the ill-posedness in the sense of Hadamard [34]. In a forthcoming paper [8], we will address the linear ill-posedness of the (unregularised) EVP model in all Sobolev spaces in a rigorous manner using this new instability (ill-posedness) that we have discussed above.
@@
We emphasise that the instability described above is fundamentally different than the (formal) ill-posedness/instability results that have been reported in the sea-ice modelling/physics literature [65, 32, 30, 31]. These results rely crucially on the absence of the viscosity cutoff (i.e. $\epsilon=0$) and are found mainly for the Hibler model (with the exception of [56, 71], which consider the (revised) EVP model as well). The rigorous analysis in [57] demonstrates that the linear ill-posedness is not present in the case of the Hibler model with $\epsilon>0$. The linear instability of the EVP model described in this section is present regardless of the value of $\epsilon$ and therefore pertains to a new kind of instability that has not been reported in the literature before, and in particular it does not apply to the Hibler model.
@@ 1.3 The main result of this paper
We consider the following regularisation of the EVP model (i.e. equations (1.3a) and (1.3c)):
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (1.10a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4\overline{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{\overline{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{\overline{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (1.10b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (1.10c)
In this model we have included the Voigt regularisation on the elastic part of the rheology in the EVP model [38] by means of the inclusion of the term $\alpha^{2}\partial_{t}\Delta\sigma$ for some fixed $\alpha>0$ in equation (1.10b). The drag forces $\mathcal{T}_{a}$ and $\mathcal{T}_{w}$ are taken as in equations (1.3d) and (1.3e), respectively. Moreover, for the sake of simplicity we have chosen $m\equiv 1$. Note that throughout this paper we will assume that $\sigma(x,t)$ is a symmetric matrix, and in Lemma 2.4 we show rigorously that the stress tensor $\sigma$ remains symmetric if $\sigma_{0}$, i.e. initially, is symmetric. Next, we will simplify the strain rate $\overline{\mathcal{D}}$ from equation (1.2) by replacing it with
$\mathcal{D}\coloneqq\lvert D(u)\rvert$, (1.11)
@@
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (1.12a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (1.12b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (1.12c)
@@
Let $u_{0}\in H^{2}(\mathbb{T}^{2})$ and $\sigma_{0}\in H^{3}(\mathbb{T}^{2})$ (which is symmetric) be given. Moreover, we assume that $P>0$ is a given constant, $U_{a},U_{w},H_{0}\in C([0,T];H^{4}(\mathbb{T}^{2}))$ and we let $\theta\in\left[0,\frac{\pi}{4}\right]$. The Voigt-EVP system (1.12) has a unique global strong solution $(u,\sigma)$ on the time interval $[0,T]$ for any $T>0$. The solution $(u,\sigma)$ enjoys the following regularity
@@
    The Voigt regularisation in equation (1.12b) is in the same spirit as the original numerical regularisation of the Hibler model from [38]. Namely, we seek an inviscid regularisation of the EVP model, which is suitable for numerical computations. The advantage of the Voigt regularisation is that on the one hand, like the original EVP model, it shares the same steady state(s) with the Hibler model. On the other hand, the Voigt regularisation constitutes a relatively modest (nonviscous) modification of the (artificial) elastic component of the rheology in the EVP model.
@@
    For the momentum equation of the Voigt-EVP model, we have omitted the advection term, in which we follow the approach in [38]. As it has been remarked in [38], the advection term is at least one order of magnitude smaller than the acceleration term and therefore it has not been included in the EVP model. In the evolution equation for the stress tensor we also do not include an advection term, as it would affect the steady states of the EVP model, which would then lose its purpose as an approximation of the Hibler model.
@@
    We note that the assumption $\theta\in\left[0,\frac{\pi}{4}\right]$ on the water turning angle, lies within the standard range in sea-ice modelling (see for example [35, 53]).